OpenStudy (anonymous):

Given cosθ=–1/5 and tan θ > 0, find sin θ.

OpenStudy (anonymous):

as tan theta >0, sin theta must be less than 0 so cos^2 +sin^2 =1 sin^2 theta=1-1/25= 24/2 \[\sin \theta = 2\sqrt6/5\]

OpenStudy (anonymous):

sorry, it will be negative

OpenStudy (anonymous):

\[-2\sqrt{6}/5\]

OpenStudy (imstuck):

OMG you poor guy. I drew it out to make it easier. Tangent is positive in quadrants 1 and 3.

OpenStudy (imstuck):

If your cosine is negative, it is negative in 2 and 3. The common quadrant here is 3 so the angle exists in Q3.

OpenStudy (imstuck):

What is the sin of that angle? Sin is opposite over hypotenuse. The side opposite is -sqrt24 and the hypotenuse is 5. So \[\sin \theta=\frac{ -\sqrt{24} }{ 5 }\]

OpenStudy (imstuck):

See that? Draw it out. It makes it so much easier.

OpenStudy (anonymous):

thank you!!

OpenStudy (imstuck):

Arnab09 did it the same way, but if you struggle with this, I find it easier to draw out the triangle and in the correct quadrant. At least that is the way I had to teach my Trig kids how to figure these out!